OpenStudy (anonymous):

@vocaloid can you help me with my hw please? c:

OpenStudy (anonymous):

I can help :)

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Ok ^-^

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

brb

OpenStudy (anonymous):

k

Vocaloid (vocaloid):

ah, I'm about to go to sleep, would you mind taking this one, _greatmath7?

OpenStudy (anonymous):

yeah @Vocaloid

OpenStudy (anonymous):

i will be glad to help

Vocaloid (vocaloid):

ty so much, gl ^ ^

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

@music.1.drug where u at

OpenStudy (anonymous):

Jabri is thinking of three numbers. The greatest number is twice as large as the least number. The middle number is three more than the least number. The sum of the three numbers is 75. Find the numbers..:P I could use the 5d process

OpenStudy (anonymous):

x + x+3 + 2x = 75 4x = 72 x=18 18, 21, 36 are the three numbers

OpenStudy (anonymous):

or G+M+L= 75 G=2L M= 3+L (2L) + (3+L) + L = 75 2L + 3 + 2L = 75 4L = 75 -3 4L = 72 L=18 G+2L = 2 *8 = 36 The numbers are -G-reatest: 36 -M-iddle: 21 -L-east: 18

OpenStudy (anonymous):

O-O that was fast.

OpenStudy (anonymous):

caculator brain

OpenStudy (anonymous):

well I have answered this question on Yahoo so it was easy I had it when I was in 6th grade

OpenStudy (anonymous):

Brb sorry

OpenStudy (anonymous):

k

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

thnx @babykel

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

welcome @_greatmath7

OpenStudy (anonymous):

:) ;0

OpenStudy (anonymous):

lol?

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

thnx @Vocaloid

OpenStudy (anonymous):

where u at? @music.1.drug

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

do u need any more help?

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

I am answering questions

OpenStudy (anonymous):

Sorry the page was freezing... and um I need help with another one ^^

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

sure